# Metastability in the Dissipative Quantum Rabi Model

### v1

@@ Abstract
The dissipative quantum Rabi model exhibits rich non-equilibrium physics, including a dissipative phase transition from the normal phase to the superradiant phase. In this work, we investigate the stability of the superradiant phase in the presence of a weak spin relaxation. We find that even a weak spin relaxation can render the superradiant phase to a superradiant metastable phase, in which symmetry-breaking states are stable only for a finite time. This arises because each spin-jump induced by relaxation applies as a strong perturbation to the system, potentially driving the system from a symmetry-breaking state to the symmetry-preserving saddle point with finite probability, before it eventually relaxes back to a symmetry-breaking state. Such dynamical processes lead to a finite lifetime of the symmetry-breaking states and restore the symmetry in the steady state. To substantiate these results, we combine mean-field and cumulant expansion analyses with exact numerical simulations. The lifetime of the symmetry-breaking states are analyzed in finite-size systems, and the conclusions are extrapolated to the thermodynamic limit via finite-size scaling. Our findings establish the metastable nature of the symmetry-breaking states in the dissipative quantum Rabi model and reveal the complexity of the dissipative phase transition beyond their equilibrium counterpart. The mechanisms uncovered here can be generalized to a broad class of open quantum systems, highlighting fundamental distinctions between equilibrium phase transitions and steady-state phase transitions.
@@ I Introduction
The quantum Rabi model (QRM), describing a two-level atom (or a spin-$1/2$) coupled to a single bosonic mode, is one of the simplest yet most profound models in quantum physics [Rabi1936]. It has inspired extensive fundamental studies that have reshaped our understanding of light–matter interactions [Brune1996, Zagoskin2008, Ashhab2010, Lo2015], strong-coupling physics [Niemczyk2010, Yoshihara2016, FornDiaz2019], integrability [Braak2011], and quantum phase transitions [Sachdev2011]. Notably, recent studies have revealed that the QRM exhibits a superradiant phase transition in the limit where the frequency ratio between the two-level atom and the bosonic mode tends to infinity [Ashhab2013, Bakemeier2012, Hwang2015, Liu2017, Zhu2020, Felicetti2020]. Moreover, subsequent studies have demonstrated that the QRM can also undergoes a dissipative phase transition when the bosonic mode is subject to relaxation, providing a versatile framework for exploring non-equilibrium critical phenomena [Hwang2018, Lyu2024, Li2024, Ilias2022, Yang2023]. These findings broadens our understanding of phase transitions by showing their occurrence even in systems comprising only a few interacting components, while also highlights promising applications in quantum critical sensing [Garbe2020, Chu2021, Beaulieu2025].
@@
Dissipative phase transition manifest as abrupt changes in the steady state of an open quantum system and exhibit strong analogies to equilibrium phase transitions through the correspondence between Liouvillian and Hamiltonian [Kessler2012, Minganti2018]. However, due to their non-equilibrium nature, dissipative phase transition exhibit richer and more complex phenomena than their equilibrium counterparts, including bistability [MendozaArenas2016, Sibalic2016, Melo2016, Alaeian2022], intermittency [Lee2012, Ates2012], multimodality [Malossi2014, Letscher2017], metastability [Macieszczak2016, Rose2016, LeBoite2017, Macieszczak2021, Jin2024], multicriticality [Soriente2018, Ferri2021], dynamical phase transition [Hedges2009, Lesanovsky2013], and phase transition without symmetry breaking [Hannukainen2018, Minganti2021]. Moreover, the stability of non-equilibrium phases under perturbations remains an open question [Letscher2017, Rakovszky2024]. In particular, metastable states can emerge in non-equilibrium systems due to the perturbations from both classical and quantum fluctuations, which undermine the stability of the non-equilibrium phase. Such states are characterized by a separation of time scales in the system dynamics [Macieszczak2016], with one timescale being much longer than all others—a feature commonly observed in finite-size systems undergoing first-order phase transition. The dissipative QRM provides an insightful starting point to explore this issue.
@@
Here, we investigate the stability of the superradiant phase in the dissipative QRM under a weak spin relaxation. The key idea is that, since only a single spin is involved in the QRM, spin quantum jumps induced by relaxation act as strong perturbations to the system and thereby undermine the stability of different phases in the dissipative QRM, particularly the symmetry-breaking phase. On the other hand, the phase diagram of the dissipative QRM is usually determined using mean-field theory. However, mean-field results become invalid when large classical or quantum fluctuations presences. For instance, multiple stationary states predicted by mean-field theory can be ruled out by more sophisticated approaches, such as variational methods [Weimer2015, Weimer2015a] and infinite tensor network simulations [Gangat2017]. Therefore, it is crucial to study the stability of the superradiant phase in dissipative QRM beyond mean-field theory. Yet, this problem has rarely been addressed in previous works.
@@
In this manuscript, we show that the superradiant phase in the dissipative QRM is generally metastable under a weak spin relaxation. This conclusion is supported by analyses based on the cumulant expansion calculations, exact Husimi $Q$-representation of the steady state, exact quantum trajectory simulations, and Liouvillian gap analysis. In particular, finite-size effects are excluded through finite-size scaling analysis. These results demonstrate that even a weak spin relaxation transforms the superradiant phase into a superradiant metastable phase, where symmetry-breaking states only has a finite lifetime. Our work provides a comprehensive analytical and numerical framework for investigating dissipative phase transitions in open quantum systems.
@@
This manuscript is organized as follows. In Sect. II, we introduce the dissipative QRM and derive its phase diagram using mean-field theory, linearly stability analysis, and cumulant expansion method. Motivated by the divergence of cumulant expansion in the superradiant phase, we analyze its stability under strong perturbations, revealing that this phase is metastable. In Sect. III, we employ the exact numerical simulations to uncover the mechanism of the metastability. The Husimi $Q$-representation of the steady-state reveals that spin relaxation opens transition channels between symmetry-breaking states, thereby leading to a finite lifetime of the symmetry-breaking states—an explicit signature of metastability. These findings are further substantiated by quantum trajectory simulations. In Sect. IV, we use the Liouvillian gap to quantify the lifetime of the symmetry-breaking states. A finite-size scaling analysis demonstrates that this lifetime remains finite in the thermodynamic limit, thereby confirming the metastability of the superradiant phase in the presence of a weak spin relaxation. Finally, we present the discussions and conclusions in Sect. V and Sect. VI, respectively.
@@ II Dissipative phase transition and stability analysis
The quantum Rabi model (QRM) describes a two-level atom (or a spin-1/2) coupled to a single bosonic mode [Braak2011, Hwang2015]. The Hamiltonian reads (with $\hbar=1$)
@@
where $\hat{a}$ ($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the bosonic mode with frequency $\omega_{0}$, $\hat{\sigma}_{x/y/z}$ are the Pauli operators of the spin with energy splitting $\Omega$, and $\lambda$ denotes the coupling strength between the spin and the bosonic mode. The QRM possesses parity symmetry, noted as $\hat{\mathcal{P}},\hat{H}]=0$ with $\hat{\mathcal{P}}\equiv e^{i\pi(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}\hat{\sigma}_{z})}$ being the parity operator. By introducing the frequency ratio $\Omega/\omega_{0}$ as a scaling parameter in the QRM, Hwang et al. revealed that the system exhibits a second-order quantum phase transition at the critical coupling strength $\lambda_{c}=\sqrt{\omega_{0}\Omega/2}$ in the limit $\Omega/\omega_{0}\rightarrow\infty$. As the coupling strength $\lambda$ changes, the ground state undergoes a transition from a parity-preserving normal phase to a parity-breaking superradiant phase [Bakemeier2012, Hwang2015, Liu2017]. Taking into account relaxation processes from both the bosonic mode and the spin due to interactions with external environments, the dissipative QRM is described by the following master equation
@@
where $\rho$ denotes the state of the system and $\mathcal{D}_{\hat{a}}\left[\rho\right]\equiv 2\kappa\hat{a}\rho\hat{a}^{\dagger}-\kappa\hat{a}^{\dagger}\hat{a}\rho-\kappa\rho\hat{a}^{\dagger}\hat{a}$ and $\mathcal{D}_{\hat{\sigma}_{-}}\left[\rho\right]\equiv 2\gamma\hat{\sigma}_{-}\rho\hat{\sigma}_{+}-\gamma\hat{\sigma}_{+}\hat{\sigma}_{-}\rho-\gamma\rho\hat{\sigma}_{+}\hat{\sigma}_{-}$ are the dissipators of the bosonic mode and the spin, with relaxation rates $\kappa$ and $\gamma$, respectively. The schematic diagram of this model is shown in Fig. 1(a). Notably, parity is still preserved under these dissipative processes, as $\hat{\mathcal{P}}(\mathcal{D}_{\hat{a}/\hat{\sigma}_{-}}[\rho])\hat{\mathcal{P}}^{\dagger}=\mathcal{D}_{\hat{a}/\hat{\sigma}_{-}}[\hat{\mathcal{P}}\rho\hat{\mathcal{P}}^{\dagger}]$. In the absence of the spin relaxation (i.e., $\gamma=0$), Hwang et al. also showed that the system still exhibits a dissipative phase transition at the critical coupling strength $\lambda_{c}=\sqrt{\frac{\omega_{0}\Omega}{2}}\sqrt{1+\frac{\kappa^{2}}{\omega_{0}^{2}}}$, where the steady state changes from the normal phase ($\lambda<\lambda_{c}$) to the superradiant phase ($\lambda>\lambda_{c}$)  [Hwang2018]. Beyond the $\gamma=0$ regime, the phase diagram and the stability of different phases, to the best of our knowledge, have received little attention.
Figure 1: (a) Schematic diagram of the dissipative quantum Rabi model under a weak spin relaxation and (b) its corresponding phase diagram. Here, NP denotes the normal phase; SP denotes the superradiant phase; and SMP denotes the superradiant metastable phase. When $\gamma=0$, the system undergoes a phase transition from the NP to the SP at the critical point $\lambda_{c}$. When $\gamma\neq 0$, this critical point extends into a critical line, and the SP transforms into the SMP. The inset shows the typical Husimi $Q$-representation of the bosonic mode in different phases.
@@ II.1 Mean-field theory
In the presence of a weak spin relaxation ($\gamma\ll\omega_{0}$), the phase diagram can be analyzed with mean-field theory. The evolution equations of observables $\left\langle\hat{x}\right\rangle\equiv\langle\frac{\hat{a}^{\dagger}+\hat{a}}{\sqrt{2}}\rangle$, $\langle\hat{p}\rangle\equiv\langle\frac{\hat{a}-\hat{a}^{\dagger}}{\sqrt{2}i}\rangle$, $\langle\hat{\sigma}_{x}\rangle$, $\langle\hat{\sigma}_{y}\rangle$, and $\langle\hat{\sigma}_{z}\rangle$ are derived from Eq. (2) by neglecting the quantum fluctuations, i.e. high-order correlations such as $\left\langle\hat{x}\hat{\sigma}_{y}\right\rangle-\langle\hat{x}\rangle\langle\hat{\sigma}_{y}\rangle$. The resulting mean-field equations read as
@@
Solving the above equations yields three steady-state solutions: one parity-preserving and a pair of parity-breaking ones. They are
@@ II.2 Linear stability analysis
Figure 2: High order cumulants $|\langle\hat{x}^{n}\rangle_{c}|$ obtained from solving the steady-state equations of cumulants truncated at 66-th order. Odd order cumulants vanish in the NP as $\langle\hat{x}^{n}\rangle_{c}=0$ holds due to the parity symmetry. Parameters are $\omega_{0}=1.0$, $\kappa/\omega_{0}=0.5$, $\gamma/\omega_{0}=0.05$, $\Omega/\omega_{0}=1200$. For NP case, $\lambda=0.6\sqrt{{\omega_{0}\Omega}/{2}}$, and for SMP case, $\lambda=1.4\sqrt{{\omega_{0}\Omega}/{2}}$.
The above linear stability analysis is applicable only when quantum fluctuations are small. However, in many open quantum systems, quantum fluctuations can be significant and may drive the system away from the steady state predicted by mean-field theory. To account for this effect, we extend the mean-field equations by incorporating quantum fluctuations through higher-order cumulants [Barquilla2020, Leymann2014, plankensteiner2022]. Formally, cumulants are defined via the generating function $\langle e^{\sum_{l}\eta_{l}\hat{o}_{l}}\rangle$ as
@@
where $n\equiv n_{1}+n_{2}+\cdots$ denotes the order of the cumulant, and $\hat{o}_{i}$ can be any operator of the bosonic mode or the spin [Kubo1962]. Higher-order cumulants ($n\geq 2$) capture deviations from mean-field theory due to quantum fluctuations. From Eq. (2), one can derive a hierarchy of equations for these cumulants, where the lower order cumulants couple with the higher order ones. To make the equations tractable, we truncate the hierarchy at a given order by neglecting all higher-order cumulants. See Appendix A for details.
In Fig. 2, we present the steady state value of $\langle\hat{x}^{n}\rangle_{c}$ obtained by solving the steady-state equations of cumulants up to $6$-th order around the the mean-field solution after neglecting $7$-th and higher order cumulants. In the NP, higher-order cumulants converge rapidly, confirming that higher-order fluctuations are negligible. This indicates that the linear stability analysis is reliable in this phase. In contrast, high-order cumulants in the parity-breaking solution diverge with the expansion order $n$ in the SMP. Such divergence indicates the presence of strong quantum fluctuations in the SMP, rendering the linear stability analysis unreliable in this regime. Our exact simulations further demonstrate that these quantum fluctuations, arising both from the intrinsic quantum fluctuation (which vanishes in the thermodynamic limit) and the spin relaxation process (which persists as long as $\gamma\neq 0$), leads to the metastability of the parity-breaking solutions.

### v2

@@ Abstract
The dissipative quantum Rabi model exhibits rich non-equilibrium physics, including a dissipative phase transition from the normal phase to the superradiant phase. In this work, we investigate the stability of the superradiant phase in the presence of a weak spin relaxation. We find that even a weak spin relaxation can render the superradiant phase a superradiant metastable phase, in which symmetry-breaking states are stable only for a finite time. This arises because each spin jump induced by relaxation applies as a strong perturbation to the system, potentially driving the system from a symmetry-breaking state to the symmetry-preserving saddle point with finite probability, before it eventually relaxes back to a symmetry-breaking state. Such dynamical processes lead to a finite lifetime of the symmetry-breaking states and restore the symmetry in the steady state. To substantiate these results, we combine mean-field and cumulant expansion analyses with exact numerical simulations. The lifetimes of the symmetry-breaking states are analyzed in finite-size systems and the conclusions are extrapolated to the thermodynamic limit via finite-size scaling. Our findings establish the metastable nature of the symmetry-breaking states in the dissipative quantum Rabi model and reveal the complexity of the dissipative phase transition beyond their equilibrium counterpart. The mechanisms uncovered here can be generalized to a broad class of open quantum systems, highlighting fundamental distinctions between equilibrium phase transitions and steady-state phase transitions.
@@ I Introduction
The quantum Rabi model (QRM), describing a two-level atom (or a spin $1/2$) coupled to a single bosonic mode, is one of the simplest yet most profound models in quantum physics [57]. It has inspired extensive fundamental studies that have reshaped our understanding of light matter interactions [10, 70, 3, 42], strong-coupling physics [53, 69, 21], integrability [8], and quantum phase transitions [61]. Notably, recent studies have revealed that the QRM exhibits a superradiant phase transition in the limit where the frequency ratio between the two-level atom and the bosonic mode tends to infinity [4, 6, 28, 41, 71, 19]. Moreover, subsequent studies have demonstrated that the QRM can also undergoes a dissipative phase transition when the bosonic mode is subject to relaxation, providing a versatile framework for exploring nonequilibrium critical phenomena [29, 44, 39, 30, 68]. These findings broaden our understanding of phase transitions by showing their occurrence even in systems comprising only a few interacting components and also highlight promising applications in quantum critical sensing [23, 13, 7].
@@
Dissipative phase transition manifest as abrupt changes in the steady state of an open quantum system and exhibit strong analogies to equilibrium phase transitions through the correspondence between the Liouvillian and Hamiltonian [32, 51]. However, due to their nonequilibrium nature, dissipative phase transitions exhibit richer and more complex phenomena than their equilibrium counterparts, including bistability [48, 63, 15, 2], intermittency [35, 5], multimodality [47, 37], metastability [45, 59, 34, 46, 31], multicriticality [64, 20], dynamical phase transitions [27, 36], and phase transition without symmetry breaking [26, 50]. Moreover, the stability of nonequilibrium phases under perturbations remains an open question [37, 58]. In particular, metastable states can emerge in nonequilibrium systems due to the perturbations from both classical and quantum fluctuations, which undermine the stability of the non-equilibrium phase. Such states are characterized by a separation of timescales in the system dynamics [45], with one timescale being much longer than all others, a feature commonly observed in finite-size systems undergoing first-order phase transition. The dissipative QRM provides an insightful starting point to explore this issue.
@@
Here we investigate the stability of the superradiant phase in the dissipative QRM under a weak spin relaxation. The key idea is that, since only a single spin is involved in the QRM, spin quantum jumps induced by relaxation act as strong perturbations to the system and thereby undermine the stability of different phases in the dissipative QRM, particularly the symmetry-breaking phase. On the other hand, the phase diagram of the dissipative QRM is usually determined using mean-field theory. However, mean-field results become invalid when large classical or quantum fluctuations are present. For instance, multiple stationary states predicted by mean-field theory can be ruled out by more sophisticated approaches, such as variational methods [66, 65] and infinite tensor network simulations [22]. Therefore, it is crucial to study the stability of the superradiant phase in the dissipative QRM beyond mean-field theory. Yet, this problem has rarely been addressed in previous works.
@@
In this paper we show that the superradiant phase in the dissipative QRM is generally metastable under a weak spin relaxation. This conclusion is supported by analyses based on the cumulant expansion calculations, exact Husimi $Q$ representation of the steady state, exact quantum trajectory simulations, and Liouvillian gap analysis. In particular, finite-size effects are excluded through finite-size-scaling analysis. These results demonstrate that even a weak spin relaxation transforms the superradiant phase into a superradiant metastable phase, where symmetry-breaking states only have a finite lifetime. Our work provides a comprehensive analytical and numerical framework for investigating dissipative phase transitions in open quantum systems.
@@
This paper is organized as follows. In Sec. II we introduce the dissipative QRM and derive its phase diagram using mean-field theory, linear stability analysis, and the cumulant expansion method. Motivated by the divergence of cumulant expansion in the superradiant phase, we analyze its stability under strong perturbations, revealing that this phase is metastable. In Sec. III we employ exact numerical simulations to uncover the mechanism of the metastability. The Husimi $Q$ representation of the steady state reveals that spin relaxation opens transition channels between symmetry-breaking states, thereby leading to a finite lifetime of the symmetry-breaking states—an explicit signature of metastability. These findings are further substantiated by quantum trajectory simulations. In Sec. IV we use the Liouvillian gap to quantify the lifetime of the symmetry-breaking states. A finite-size-scaling analysis demonstrates that this lifetime remains finite in the thermodynamic limit, thereby confirming the metastability of the superradiant phase in the presence of a weak spin relaxation. We present a discussion and conclusion in Sec. V and summarize in Sect. VI.
@@ II Dissipative phase transition and stability analysis
The quantum Rabi model describes a two-level atom (or a spin-1/2) coupled to a single bosonic mode [8, 28]. The Hamiltonian reads (with $\hbar=1$)
@@
where $\hat{a}$ ($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the bosonic mode with frequency $\omega_{0}$, $\hat{\sigma}_{x,y,z}$ are the Pauli operators of the spin with energy splitting $\Omega$, and $\lambda$ denotes the coupling strength between the spin and the bosonic mode. The QRM possesses parity symmetry, denoted by $\hat{\mathcal{P}},\hat{H}]=0$ with $\hat{\mathcal{P}}\equiv e^{i\pi(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}\hat{\sigma}_{z})}$ the parity operator. By introducing the frequency ratio $\Omega/\omega_{0}$ as a scaling parameter in the QRM, Hwang et al. revealed that the system exhibits a second-order quantum phase transition at the critical coupling strength $\lambda_{c}=\sqrt{\omega_{0}\Omega/2}$ in the limit $\Omega/\omega_{0}\rightarrow\infty$. As the coupling strength $\lambda$ changes, the ground state undergoes a transition from a parity-preserving normal phase to a parity-breaking superradiant phase [6, 28, 41]. Taking into account relaxation processes from both the bosonic mode and the spin due to interactions with external environments, the dissipative QRM is described by the master equation
@@
where $\rho$ denotes the state of the system and $\mathcal{D}_{\hat{a}}\left[\rho\right]\equiv 2\kappa\hat{a}\rho\hat{a}^{\dagger}-\kappa\hat{a}^{\dagger}\hat{a}\rho-\kappa\rho\hat{a}^{\dagger}\hat{a}$ and $\mathcal{D}_{\hat{\sigma}_{-}}\left[\rho\right]\equiv 2\gamma\hat{\sigma}_{-}\rho\hat{\sigma}_{+}-\gamma\hat{\sigma}_{+}\hat{\sigma}_{-}\rho-\gamma\rho\hat{\sigma}_{+}\hat{\sigma}_{-}$ are the dissipators of the bosonic mode and the spin, with relaxation rates $\kappa$ and $\gamma$, respectively. The schematic diagram of this model is shown in Fig. 1(a). Notably, parity is still preserved under these dissipative processes, as $\hat{\mathcal{P}}(\mathcal{D}_{\hat{a}/\hat{\sigma}_{-}}[\rho])\hat{\mathcal{P}}^{\dagger}=\mathcal{D}_{\hat{a}/\hat{\sigma}_{-}}[\hat{\mathcal{P}}\rho\hat{\mathcal{P}}^{\dagger}]$. In the absence of spin relaxation, i.e., $\gamma=0$, Hwang et al. also showed that the system still exhibits a dissipative phase transition at the critical coupling strength $\lambda_{c}=\sqrt{\frac{\omega_{0}\Omega}{2}}\sqrt{1+\frac{\kappa^{2}}{\omega_{0}^{2}}}$, where the steady state changes from the normal phase ($\lambda<\lambda_{c}$) to the superradiant phase ($\lambda>\lambda_{c}$)  [29]. Beyond the $\gamma=0$ regime, the phase diagram and the stability of different phases have received little attention.
Figure 1: (a) Schematic diagram of the dissipative quantum Rabi model under a weak spin relaxation and (b) its corresponding phase diagram. Here, NP denotes the normal phase, SP denotes the superradiant phase, and SMP denotes the superradiant metastable phase. When $\gamma=0$, the system undergoes a phase transition from the NP to the SP at the critical point $\lambda_{c}$. When $\gamma\neq 0$, this critical point extends into a critical line, and the SP transforms into the SMP. The inset shows the typical Husimi $Q$ representation of the bosonic mode in different phases.
@@ II.1 Mean-field theory
In the presence of a weak spin relaxation ($\gamma\ll\omega_{0}$), the phase diagram can be analyzed with mean-field theory. The evolution equations of observables $\left\langle\hat{x}\right\rangle\equiv\langle\frac{\hat{a}^{\dagger}+\hat{a}}{\sqrt{2}}\rangle$, $\langle\hat{p}\rangle\equiv\langle\frac{\hat{a}-\hat{a}^{\dagger}}{\sqrt{2}i}\rangle$, $\langle\hat{\sigma}_{x}\rangle$, $\langle\hat{\sigma}_{y}\rangle$, and $\langle\hat{\sigma}_{z}\rangle$ are derived from Eq. (2) by neglecting the quantum fluctuations, i.e. high-order correlations such as $\left\langle\hat{x}\hat{\sigma}_{y}\right\rangle-\langle\hat{x}\rangle\langle\hat{\sigma}_{y}\rangle$. The resulting mean-field equations are
@@
Solving the above equations yields three steady-state solutions: one parity-preserving and a pair of parity-breaking ones
@@ II.2 Linear stability analysis
Figure 2: High-order cumulants $|\langle\hat{x}^{n}\rangle_{c}|$ obtained from solving the steady-state equations of cumulants truncated at sixth order. Odd order cumulants vanish in the NP as $\langle\hat{x}^{n}\rangle_{c}=0$ holds due to the parity symmetry. The parameters are $\omega_{0}=1.0$, $\kappa/\omega_{0}=0.5$, $\gamma/\omega_{0}=0.05$, and $\Omega/\omega_{0}=1200$. For the NP case, $\lambda=0.6\sqrt{{\omega_{0}\Omega}/{2}}$, and for the SMP case, $\lambda=1.4\sqrt{{\omega_{0}\Omega}/{2}}$.
The above linear stability analysis is applicable only when quantum fluctuations are small. However, in many open quantum systems, quantum fluctuations can be significant and may drive the system away from the steady state predicted by mean-field theory. To account for this effect, we extend the mean-field equations by incorporating quantum fluctuations through higher-order cumulants [62, 38, 54]. Formally, cumulants are defined via the generating function $\langle e^{\sum_{l}\eta_{l}\hat{o}_{l}}\rangle$ as
@@
where $n\equiv n_{1}+n_{2}+\cdots$ denotes the order of the cumulant and $\hat{o}_{i}$ can be any operator of the bosonic mode or the spin [1]. Higher-order cumulants ($n\geq 2$) capture deviations from mean-field theory due to quantum fluctuations. From Eq. (2) one can derive a hierarchy of equations for these cumulants, where the lower-order cumulants couple with the higher-order ones. To make the equations tractable, we truncate the hierarchy at a given order by neglecting all higher-order cumulants (see Appendix A for details).
In Fig. 2 we present the steady state value of $\langle\hat{x}^{n}\rangle_{c}$ obtained by solving the steady-state equations of cumulants up to sixth order around the mean-field solution after neglecting seventh and higher-order cumulants. In the NP, higher-order cumulants converge rapidly, confirming that higher-order fluctuations are negligible. This indicates that the linear stability analysis is reliable in this phase. In contrast, high-order cumulants in the parity-breaking solution diverge with the expansion order $n$ in the SMP. Such divergence indicates the presence of strong quantum fluctuations in the SMP, rendering the linear stability analysis unreliable in this regime. Our exact simulations further demonstrate that these quantum fluctuations, arising from both the intrinsic quantum fluctuation (which vanishes in the thermodynamic limit) and the spin relaxation process (which persists as long as $\gamma\neq 0$), lead to the metastability of the parity-breaking solutions.